# Video Soundtrack Generation by Aligning Emotions and Temporal Boundaries

### v2

@@ 4.3 Objective evaluation
We first evaluate music similarity using the ground-truth audio from the EmoMV-C dataset. Although the ground truth is not in MIDI, all methods generate MIDI outputs that are synthesized to audio, enabling a fair comparison in the audio domain. We evaluate the Kullback-Leibler Divergence (KL) between the labels of generated and ground-truth music, calculated using a pretrained music tagging model (Won et al. 2020). This measures how closely the distribution of semantic tags in the generated music matches that of the ground-truth. We avoid using Fréchet Audio Distance, as it compares overall audio distributions suited for unconditional generation (Kilgour et al. 2019), whereas video-based music generation requires per-sample evaluation to assess alignment with each video’s unique content.
@@
We additionally evaluate temporal audio-video alignment (AV-Align) score, which is the Intersection-over-Union between video and audio peaks (Yariv et al. 2024), using both EmoMV-C and Ads datasets. Video peaks are detected from the mean optical flow magnitude (Horn and Schunck 1981), while audio peaks correspond to audio onsets (Böck and Widmer 2013). We report AV-Align scores using windows of one video frame (33.4 ms at 30 FPS) and one second. Each window is applied to detect local peaks in both video and audio and then used to match peaks within the same temporal range for alignment.
4.4 Subjective evaluation
@@
Using the evaluation videos from the EmoMV-C and the Ads datasets, we generate accompanying music with the three models and create survey pages, each containing two videos. For each video, the three music versions generated by the compared models are presented side by side with anonymized model names, with each model’s output appearing equally in the left, center, and right positions. We also provide output samples from eight videos online, with model names deanonymized and each model’s output appearing in the same position for clarity. [^3] https://em-sync.github.io
We enrolled 153 participants and asked them to rank the three models based on the standard criteria used in previous works (Di et al. 2021; Kang, Poria, and Herremans 2024): Music Richness (MR) — the richness and diversity of the music, independent of the video content; Music Quality (MQ) — the overall quality of the music, independent of the video content; Emotion Match (EM) — how well the music matches the video in terms of emotion; Timing Match (TM) — how well the music synchronizes with the video in terms of rhythm and timing; and Overall Match (OM) — how well the music matches the video as a whole. Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst).
5 Results and conclusion
Table 2 presents objective evaluation results for EMSYNC, Video2Music (Kang, Poria, and Herremans 2024), and CMT (Di et al. 2021), on the EmoMV-C and Ads datasets. EMSYNC outperforms both baselines across a majority of the metrics. On EmoMV-C, it achieves the best frame-level alignment and the lowest KL divergence, indicating superior synchronization to video and similarity to ground-truth music. While Video2Music performs best on second-level alignment, EMSYNC remains competitive. On the Ads dataset, EMSYNC obtains the best scores across both alignment metrics, further demonstrating its robustness across different video domains. We don’t calculate KL divergence because the Ads dataset doesn’t include ground truth music.
Table 3 demonstrates the subjective results in terms of average user rankings with 95% confidence margins. A lower average score represents a better ranking, with 1 being the best and 3 the worst possible. EMSYNC achieves the best rankings across all metrics and datasets. The difference is more pronounced in video-music correspondence metrics such as emotion match and timing match, highlighting our model’s stronger alignment to video. For overall match, our method outperforms the others with non-overlapping 95% confidence intervals, reflecting a decisive user preference for EMSYNC in video-based music generation.
These results highlight EMSYNC’s effectiveness in generating music that aligns with the input video both temporally and emotionally, across objective and subjective evaluations. By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators.
@@ 5 Results and conclusion
Metric AVf AVs KL AVf AVs
CMT 0.014 0.150 4.728 0.008 0.099
Video2Music 0.010 0.245 9.789 0.002 0.107
EMSYNC (Ours) 0.017 0.233 4.175 0.011 0.150
Table 2: Objective results: AVf and AVs denote audio-video alignment in frame and second-levels respectively, where higher is better. KL denotes KL-divergence where lower is better.
D M CMT Video2Music
EMSYNC
EmoMV-C
MR 1.870 ± 0.102 2.317 ± 0.108 1.812 ± 0.112
@@ D M CMT Video2Music
Table 3: Subjective results: Mean ranks ± 95% confidence margins for the subjective metrics, where 1 is the best possible and 3 is the worst possible ranking. The first two columns represent the dataset and metric, indicated by the headers D and M. The table is transposed for compactness.

### v3

@@ IV-C Objective evaluation
We first evaluate music similarity using the ground-truth audio from the EmoMV-C dataset. Although the ground truth is not in MIDI, all methods generate MIDI outputs that are synthesized to audio, enabling a fair comparison in the audio domain. We evaluate the Kullback-Leibler Divergence (KL) between the labels of generated and ground-truth music, calculated using a pretrained music tagging model [50]. This measures how closely the distribution of semantic tags in the generated music matches that of the ground-truth. We avoid the commonly used Fréchet Audio Distance, as it compares overall audio distributions suited for unconditional generation [23], whereas video-based music generation requires per-sample evaluation to assess alignment with each video’s unique content.
@@
We additionally evaluate temporal audio-video alignment (AV-Align) score, which is the Intersection-over-Union between video and audio peaks [53], using both EmoMV-C and Ads datasets. Video peaks are detected from the mean optical flow magnitude [16], while audio peaks correspond to audio onsets [3]. We calculate AV-Align scores at two temporal resolutions (windows): one video frame (33.3 ms at 30 FPS) and one second. For each window, local peaks in both video and audio are detected and matched within the same temporal range to assess alignment. As an internal evaluation, we additionally report the alignment between the generated MIDI chords (specifically the CHORD token) and video scene cuts using the same 33.3 ms and 1 s windows. As baseline methods lack an explicit boundary-matching architecture, these alignment scores are reported exclusively for EMSYNC to validate the effectiveness of its temporal conditioning mechanism. Finally, we report the average runtime for each model when generating music for a one-minute video, including model initialization, video analysis, and music generation.
IV-D Subjective evaluation
@@
Using the evaluation videos from the EmoMV-C and the Ads datasets, we generate accompanying music with the three models and create survey pages, each containing two videos. For each video, the three music versions generated by the compared models are presented side by side with anonymized model names, with each model’s output appearing an equal number of times in the left, center, and right positions. We also provide output samples from eight videos online, with model names deanonymized and each model’s output appearing in the same position for clarity. [^3] https://serkansulun.com/emsync
We enrolled 153 remote participants who provided informed consent before taking part in the survey. They were asked to rank the three models based on the standard criteria used in previous works [8, 21]:
  • Music Richness (MR) — the richness and diversity of the music, independent of the video content.
  • Music Quality (MQ) — the overall quality of the music, independent of the video content.
  • Emotion Match (EM) — how well the music matches the video in terms of emotion.
  • Timing Match (TM) — how well the music synchronizes with the video in terms of rhythm and timing.
  • Overall Match (OM) — how well the music matches the video as a whole.
Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst). Figure 4 shows a sample view of the user survey.
V Results and conclusion
TABLE II: Objective comparison with the state-of-the-art
@@ V Results and conclusion
Metric AVf AVs KL AVf AVs RT
CMT 0.014 0.150 4.728 0.008 0.099 3.55
Video2Music 0.010 0.245 9.789 0.002 0.107 1.61
EMSYNC (Ours) 0.017 0.233 4.175 0.011 0.150 1.42
Table II presents objective evaluation results for EMSYNC, Video2Music [21], and CMT [8], on the EmoMV-C and Ads datasets. AVf and AVs denote audio-video alignment in frame and second-levels, respectively, where higher is better. KL denotes the KL-divergence, where lower is better. RT denotes average runtime in minutes to generate music for a 1-minute video, including model initialization, video analysis, and music generation, where lower is better. EMSYNC outperforms both baselines across a majority of the metrics. On EmoMV-C, it achieves the best frame-level alignment and the lowest KL divergence, indicating superior synchronization to video and similarity to ground-truth music. While Video2Music performs best on second-level alignment, EMSYNC remains competitive. On the Ads dataset, EMSYNC obtains the best scores across both alignment metrics, further demonstrating its robustness across different video domains. We don’t calculate KL divergence because the Ads dataset doesn’t include ground truth music. EMSYNC also achieves the fastest runtime among the evaluated methods. For the EmoMV-C and Ads datasets, EMSYNC achieves chord-scene cut alignment scores of 0.297 and 0.317 at the 33.3 ms resolution, and 0.899 and 0.909 within a one-second window, respectively, demonstrating its capabilities in aligning temporal boundaries through its temporal conditioning mechanism.
TABLE III: Subjective comparison with the state-of-the-art:
Mean ranks ± 95% confidence margins
Dataset Metric CMT Video2Music EMSYNC (Ours)
EmoMV-C MR 1.870 ± 0.102 2.317 ± 0.108 1.812 ± 0.112
@@
Table III demonstrates the subjective results in terms of average user rankings with 95% confidence margins. A lower average score represents a better ranking, with 1 being the best and 3 the worst possible. EMSYNC achieves the best average ranking across all metrics and datasets. The difference is more pronounced in video-music correspondence metrics such as emotion match and timing match, highlighting our model’s stronger alignment to video. For the overall match, our method outperforms the others with non-overlapping 95% confidence intervals, reflecting a decisive user preference for EMSYNC in video-based music generation.
Figure 5 further shows the distribution of rankings assigned by participants in the subjective evaluation. EMSYNC achieves the highest frequency of the top rank (Rank 1) across all metrics and datasets.
Figure 5: Distributions of rankings assigned by the participants in the subjective evaluation using EmoMV-C (top) and Ads (bottom) datasets.
These results highlight EMSYNC’s effectiveness in generating music that aligns with the input video both temporally and emotionally, across objective and subjective evaluations. By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators. In future work, we aim to incorporate more complex musical boundaries, such as those that separate melodic segments or motifs. We will also enable dynamic emotional conditioning to adapt the music continuously to evolving complex emotional content in videos.
VI Acknowledgments
This work is co-financed by Component 5 - Capitalization and Business Innovation, integrated in the Resilience Dimension of the Recovery and Resilience Plan within the scope of the Recovery and Resilience Mechanism (MRR) of the European Union (EU), framed in the Next Generation EU, for the period 2021 - 2026, within project NEXUS, with reference 53, and the fellowship from FCT - Fundação para a Ciência e a Tecnologia (2022.09594.BD).